OpenStudy (anonymous):

math homework!!! please help create a sample problem that has a binomial which is a factor of the polynomial being divided, aand another problem that has a binomial which is not a factor of the polynomial being divided

OpenStudy (is3535):

3x^3 - 2x^2 - 2x - 5) / (3x - 5) = x^2 + x + 1 Since there is no remainder, 3x - 5 is a factor of 3x^3 - 2x^2 - 2x - 5. (3x^3 - 2x^2 - 2x - 7) / (3x - 5) = x^2 + x + 1 with remainder -2. Since the remainder is not zero, 3x - 5 is not a factor of 3x^3 - 2x^2 - 2x - 7.

OpenStudy (anonymous):

tthank you :))